OpenStudy (anonymous):

this fraction: (a^2-5)/(a^3-1) minus this fraction (a+1)/(a^2+a+1) a. (-4)/(a^3-1) b. (-6)/(a^3-1) c. (a^2-a-6)/(2a) ^^^^^^^^^^^^^^^^ all of these are fractions

OpenStudy (anonymous):

 If this helps, this is the problem

OpenStudy (campbell_st):

well you need a common denominattor the factored form is the denominator in the 1st fraction is a^3 -1 = (a -1)(a^2 + a + 1) this is the difference of 2 cubes Then multiply the numerator and denominator of the 2nd fraction by (a - 1) which would give \[\frac{a^2 - 5}{(a - 1)(a^2 + a + 1} - \frac{(a + 1)(a -1)}{(a-1)(a^2 + a + 1)}\] now simplify it by collecting like terms in the numerator

OpenStudy (campbell_st):

lookin at the answers you will need to rewrite the denominator as the difference of 2 cubes that is a^3 - 1

OpenStudy (anonymous):

Thank you @campbell_st

OpenStudy (anonymous):

I got the first answer.

OpenStudy (campbell_st):

that's what I got